A cannonball is fired with a velocity of 100m/s at 25degrees above the horizontal. Determine how far away it lands on level ground.
1
Expert's answer
2019-11-11T15:52:43-0500
"R=\frac{v^2}{g}\sin{2(25)}"

"R=\frac{100^2}{9.8}\sin{50}=782\ m"